Do one of the​ following, as appropriate.​ (a) Find the critical value z Subscript alpha divided by 2​, ​(b) find the critical value t Subscript alpha divided by 2​, ​(c) state that neither the normal nor the t distribution applies. Confidence level 95​%; nequals26​; sigma equals 30.2​; population appears to be normally distributed.

Find the critical value.

A. zα/2 = 1.645

B. zα/2 = 1.96

C. tα/2 = 2.060

D. tα/2 = 1.708

E. Neither normal nor t distribution applies.
Answer #1

SOLUTION:

given that n=26

HERE population standard deviation known sigma=30.2

so we can use z table

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

correct option B)1.96